### v1

# Resonant excitation of quasinormal modes of black holes

Abstract
We elucidate that a distinctive resonant excitation between quasinormal modes (QNMs) of black holes emerges as a universal phenomenon at avoided crossing near exceptional point through high-precision numerical analysis and theory of QNMs based on the framework of non-Hermitian physics. This resonant phenomenon not only allows us to decipher a long-standing mystery concerning the peculiar behaviors of QNMs but also stands as a novel beacon for characterizing black hole spacetime geometry. Our findings pave the way for rigorous examinations of black holes and the exploration of new physics in gravity.
Introduction. The last decade has witnessed the advent of gravitational wave (GW) astronomy, opening a new window to the universe, especially in the dynamical and strong-field regime near black holes Abbott et al. (2016); Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021). General relativity predicts that black holes are the simplest astrophysical objects in the universe since the only rotating astrophysical black hole solution that satisfies the Einstein equation is the Kerr spacetime Israel (1967); Carter (1971); Hawking (1972), characterized solely by mass and spin. Remnant black holes after binary black hole mergers are expected to settle into the state of a Kerr black hole, emitting ringdown GWs. Linear perturbation theory predicts that the ringdown GWs are described by a linear combination of intrinsic exponentially damped sinusoids known as quasinormal modes (QNMs) Vishveshwara (1970); Press (1971); Teukolsky (1973); Chandrasekhar and Detweiler (1975). Their timescales and amplitudes are governed by complex frequencies and excitation factors Leaver (1985, 1986a, 1986b); Sun and Price (1988); Nollert and Schmidt (1992); Nollert and Price (1999); Andersson (1995, 1997); Glampedakis and Andersson (2001, 2003), also determined solely by mass and spin, playing a central role in the black hole spectroscopy program Detweiler (1980); Echeverria (1989); Finn (1992); Dreyer et al. (2004); Berti et al. (2006). Despite extensive efforts to better understand them, the peculiar behaviors of QNMs reflecting changes in black hole parameters have remained elusive Kokkotas and Schmidt (1999); Nollert (1999); Berti et al. (2009); Konoplya and Zhidenko (2011).
In this Letter, we reveal the emergence of the resonant excitation phenomenon associated with an avoided crossing of QNMs near the exceptional point (EP). The avoided crossing in quantum mechanics, also known as level repulsion, is a phenomenon where two energy eigenvalues cannot coincide unless a certain condition is satisfied Hund (1927); von Neuman and Wigner (1929); Landau and Lifshitz (1977); Arnold (1978). It plays a pivotal role in experiments and observations across vast fields of physical science Ashcroft and Mermin (1976); Landau (1932); Zener (1932); Majorana (1932); Stückelberg (1932); Ivakhnenko et al. (2023); Purcell (1946); Herzberg (1991); Smirnov (2003); Wurm (2017); Giganti et al. (2018). The EP Kato (1995); Özdemir et al. (2019); Wiersig (2020); Parto et al. (2021); Ding et al. (2022), a branch point singularity in the complex eigenvalue plane, is crucial in non-Hermitian physics El-Ganainy et al. (2018); Bergholtz et al. (2021); Ashida et al. (2021), where avoided crossing occurs in its vicinity Heiss (2000, 2012). With these concepts, we formulate the resonance of QNMs and provide novel perspectives for exploring gravity. We employ geometric units with $G=c=1$ throughout.
Kerr QNM frequencies and excitation factors. We first reveal the emergence of the resonance of Kerr QNMs at avoided crossing and identify their unique features. Linear perturbation theory in general relativity predicts that the ringdown GW strain $h=h_{+}-ih_{\times}$ at far distance from perturbed Kerr black hole is given by a superposition of QNM damped sinusoids Vishveshwara (1970); Press (1971); Teukolsky (1973); Chandrasekhar and Detweiler (1975):
$h=\sum_{\ell,m,n}\frac{I_{\ell mn}B_{\ell mn}S_{\ell mn}e^{im\varphi}}{\omega_{\ell mn}^{2}}e^{-i\omega_{\ell mn}t},$ (1)
in a decomposition in terms of spin-weighted spheroidal harmonics $S_{\ell mn}(\theta,a\omega_{\ell mn})$ Teukolsky (1972, 1973); Press and Teukolsky (1973). While a complex factor $I_{\ell mn}$ involves an integral depending on the initial condition, recent studies suggest that for each $(\ell,m)$ multipole $I_{\ell mn}$ does not strongly depend on the overtone index $n$ and black hole spin Oshita (2021); Cheung et al. (2024). Therefore, the ringdown GWs are mainly governed by complex QNM frequencies $\omega_{\ell mn}$ and excitation factors $B_{\ell mn}$ defined by Leaver (1986b)
$B_{\ell mn}=\left.\frac{A_{\rm out}}{2\omega}\left(\frac{\mathop{}\!\mathrm{d}A_{\rm in}}{\mathop{}\!\mathrm{d}\omega}\right)^{-1}\right|_{\omega=\omega_{\ell mn}},$ (2)
with asymptotic amplitudes $A_{\rm out}$ and $A_{\rm in}$ for the outgoing and ingoing waves, respectively. $\omega_{\ell mn}$ and $B_{\ell mn}$ depend only on mass $M$ and spin $a$ of the black hole. Since the mass dependency is a simple scaling, the dependency on a dimensionless spin parameter $a/M$ is of primary interest.
Figure 1: Kerr QNM frequencies $\omega_{\ell mn}$ (left) of $(\ell,m)=(2,2)$ from $n=0$ (bottom red) to $7$ (top magenta) for $0\leq a/M\leq 1-10^{-6}$ with arrows indicating the direction of increasing spin. Excitation factors $B_{\ell mn}$ with $n=5,6$ (middle) and other overtones (right), the latter of which is normalized. The range $a/M=0.875$ (circle) – $0.915$ (triangle) is highlighted with small dots with the spacing of $0.005$.
Figure 2: Kerr QNM frequencies $\omega_{\ell mn}$ (top) and excitation factors $B_{\ell mn}$ (bottom) of $(\ell,m)=(3,1)$ from $n=4$ to $7$ for $0.95\leq a/M\leq 1-10^{-6}$. Highlighted are the ranges $a/M=0.952$ (diamond) – $0.96$ (square), $0.9715$ (circle) – $0.973$ (triangle), and $0.978$ (plus) – $0.983$ (cross) with small dots with the spacing of $5\times 10^{-4}$, $5\times 10^{-5}$, and $5\times 10^{-4}$, respectively. Black dashed curves represent the hyperbola (top) and lemniscate (bottom).
We revisit numerical calculation of the QNM frequencies and excitation factors of Kerr black holes. We closely follow the strategy adopted in Zhang et al. (2013) and implement some technical improvements to achieve high-precision calculation up to higher overtones which will be presented elsewhere  [^1] In addition, we also correct some errors found in an intermediate step of Zhang et al. (2013). Consequently, our results of excitation factors and the ones in Zhang et al. (2013) differ for a factor of $e^{-i\omega(1-\sqrt{1-(a/M)^{2}})}$.. The dataset is publicly available Motohashi (2024a).
For $(\ell,m)=(2,2)$ multipole of GWs, most of the Kerr QNM frequencies uniformly migrate towards the accumulation point at $2M\omega_{\ell mn}=m$ Detweiler (1980); Glampedakis and Andersson (2001); Cardoso (2004); Hod (2008); Yang et al. (2013a, b). However, as depicted in Fig. 1, only the fifth overtone suddenly turns over at $a/M\simeq 0.9$ with making a small knot-shape loop and heads towards an isolated point far from the accumulation point. This behavior was first pointed out by Onozawa Onozawa (1997) about three decades ago and has been confirmed by other work Yang et al. (2013b); Cook and Zalutskiy (2014), but the physical reason has been veiled. While it has been implicitly assumed that the fifth overtone solely behaves in a strange manner, here we point out that, while less manifest, the trajectory of the sixth overtone QNM frequency is slightly distorted when it passes near the fifth overtone at $a/M\simeq 0.9$.
On the other hand, it was recently found that the absolute values of $B_{225}$ and $B_{226}$ become large for high-spin Kerr black holes Giesler et al. (2019); Oshita (2021), despite the fact that the excitation factors tend to zero at the extremal spin Ferrari and Mashhoon (1984); Berti and Cardoso (2006); Zhang et al. (2013) and asymptotically scale as $n^{-1}$ for higher overtones Andersson (1997); Berti and Cardoso (2006). Again, the physical reason has been unclear. However, when viewed on the complex plane in Fig. 1, we notice that the excitation factors actually exhibit a suggestive behavior. While other overtones follow spiral trajectories, it turns out that $B_{225}$ and $B_{226}$ are against the spirals and are enhanced in almost opposite directions at $a/M\simeq 0.9$.
The implications of these observations are two-fold: First, this phenomenon should be viewed as occurring in a pair of two modes rather than a single mode. Second, there should be an underlying interplay between the turning/distortion of the QNM frequencies and the strong excitations.
Actually, a similar phenomenon occurs more manifestly for higher multipoles. In Fig. 2, we see that $(\ell,m)=(3,1)$ modes exhibit the distinctive behaviors successively between multiple pairs. Among them the pair of $n=4$ and $5$ is the most manifest case, exhibiting a sharp repulsion between the two QNM frequencies  [^2] Such repulsion was also observed in the QNMs of scalar field in Kerr Yang et al. (2013b), (anti-)de Sitter black holes Jansen (2017); Davey et al. (2022); Kinoshita et al. (2024), and charged black holes Dias et al. (2022a, b); Davey et al. (2024).. In Fig. 2, we also note that the mild repulsion between $n=6$ and $7$ overtones seems to be a trigger of the looping trajectory of the sixth overtone found in Onozawa (1997); Cook and Zalutskiy (2014). Further, during these peculiar behaviors of QNM frequencies, the corresponding pairs of excitation factors exhibit symmetric amplifications. The sharper the repulsion between two QNM frequencies are, the more strongly and clearly point symmetrically the corresponding pair of excitation factors are amplified.
Interestingly, for the most manifest case in Fig. 2, we find that the QNM frequencies and excitation factors almost follow the hyperbola and the lemniscate of Bernoulli, respectively. When written down in the polar coordinates with $z=re^{i\theta}$, the hyperbola $r^{-2}=\cos 2\theta$ and lemniscate $r^{2}=\cos 2\theta$ are inverses of each other. Albeit approximate, it is remarkable that such intimately related curves are inherent in QNM spectrum of Kerr GWs.
Figure 3: Absolute values of the difference between the excitation factors (blue solid) and the inverse of the difference between QNM frequencies (orange dashed), the latter of which is normalized to match the maxima, with a fit by a quarter-power Lorentzian (black dotted).
We show in Fig. 3 that the inverse proportionality between the difference of excitation factors $\Delta B$ and the difference of QNM frequencies $\Delta\omega$ is actually a common feature of the pair excitations, providing insight that they belong to the same underlying phenomenon. As expected, the inverse proportionality is more manifest for the case of sharper repulsion and stronger excitation. Further, we find that the peaks in Fig. 3, especially the sharp ones, demonstrate an excellent agreement with a quarter-power Lorentzian $f(a/M)$, where $f(x)=f_{0}[\gamma^{2}/\{(x-x_{0})^{2}+\gamma^{2}\}]^{1/4}$ with positive parameters $f_{0},\gamma,x_{0}$, rather than the Lorentzian itself. Not only the absolute value, we also confirm that a product $\Delta B\Delta\omega$ remains almost constant as a complex number during the resonant excitation.
We find that this phenomenon ubiquitously occurs for Kerr QNMs, i.e., GWs with other multipoles, and scalar and electromagnetic fields. Further, even in QNMs away from Kerr, we find that the resonances with the same features show up (see Supplemental Material). These are tantalizing hints to a universal nature of QNMs.
Theory of quasinormal modes. Below we provide a theoretical explanation for this phenomenon by developing a theory of QNMs. Let us consider the QNMs of a general master equation for perturbations $\psi$ around a black hole after Fourier and harmonic expansions,
$\left(\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}+U\right)\psi=0,$ (3)
where $U$ is an effective potential and $x$ denotes the tortoise coordinate with $x\to\pm\infty$ corresponding to the spatial infinity and event horizon of the black hole, respectively. For nonrotating black holes, $U=\omega^{2}-V(x)$ Regge and Wheeler (1957); Zerilli (1970) and hence (3) reads $H\psi=\omega^{2}\psi$ with an effective Hamiltonian $H=-\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}+V(x)$, which is analogous to the Schrödinger equation in quantum mechanics. For radial perturbation equation of rotating black holes, $U$ depends on $x,\omega$ and other parameters Teukolsky (1972, 1973), so (3) defines a nonlinear eigenvalue problem. $U$ changes as the spin parameter varies, causing the migration of QNM frequencies.
A crucial difference of QNMs from the bound state in quantum mechanics originates from the boundary condition. Instead of decaying at infinity, the QNM wave function is required to satisfy the in/outgoing wave condition $\psi\to e^{\pm i\omega x}$ as $x\to\pm\infty$. This Siegert boundary condition discretizes the QNM eigenvalues as complex number with negative imaginary part. Hence, the QNM wave function diverges at $x\to\pm\infty$ (when viewed at constant time hypersurface), and the standard inner product for bound states is not well-defined for QNMs. Several kinds of definitions of the inner product for QNMs of black holes have been considered Ching et al. (1995); Leung et al. (1997, 1998); Yang et al. (2015); Zimmerman et al. (2015); Jaramillo et al. (2021); Gasperin and Jaramillo (2022); Yang and Zhang (2023); Green et al. (2023); London (2023); London and Gurevich (2023); Ma and Yang (2024).
In quantum mechanics, this kind of system is known as the Schrödinger problem of open systems and has been utilized for studies of quantum resonances, for which the Hamiltonian is not Hermitian Gamow (1928); Landau and Lifshitz (1977); Kukulin et al. (1989); Moiseyev (2011). More recently, the non-Hermitian physics has been extensively studied both from theoretical and experimental point of view El-Ganainy et al. (2018); Bergholtz et al. (2021); Ashida et al. (2021). Here, we define a regular inner product, or more precisely biorthogonal product, for QNMs inspired by the case of the quantum resonance states Kukulin et al. (1989); More (1971a, b); Sternheim and Walker (1972). It corresponds to a generalization of Ching et al. (1995); Leung et al. (1997, 1998) to the case with (3) including rotating black holes. In this formulation, QNM wave function $\psi_{n}$ with eigenvalue $\omega_{n}$ corresponds to the resonance state, for which we introduce so-called anti-resonance (or adjoint) state $\tilde{\psi}_{n}\coloneqq\psi_{n}^{*}$ with eigenvalue $\tilde{\omega}_{n}\coloneqq-\omega_{n}^{*}$. We then define the biorthogonal product for QNMs along a contour $\mathcal{C}$ as
$(\tilde{\psi}_{n}|\psi_{k})\coloneqq\int_{\mathcal{C}}\mathop{}\!\mathrm{d}{x}\frac{U_{n}-U_{k}}{\omega_{n}^{2}-\omega_{k}^{2}}\psi_{n}\psi_{k}-\frac{[\psi_{n}\psi^{\prime}_{k}-\psi^{\prime}_{n}\psi_{k}]_{\partial\mathcal{C}}}{\omega_{n}^{2}-\omega_{k}^{2}},$ (4)
where $U_{n}=U|_{\omega=\omega_{n}}$. From (3), it is straightforward to show that the biorthogonality exactly holds for any $\mathcal{C}$, i.e., $(\tilde{\psi}_{n}|\psi_{k})=0$ when $n\neq k$. Formally taking the limit $\omega_{k}\to\omega_{n}$, we define the “norm squared” for QNMs as
$\mathcal{A}_{n}^{2}\coloneqq\int_{\mathcal{C}}\mathop{}\!\mathrm{d}{x}\frac{\partial U_{n}}{\partial\omega_{n}^{2}}\psi_{n}^{2}-\left.\frac{\partial}{\partial\omega_{k}^{2}}[\psi^{\prime}_{n}\psi_{k}-\psi_{n}\psi^{\prime}_{k}]_{\partial\mathcal{C}}\right|_{\omega_{k}=\omega_{n}},$ (5)
which is in general complex  [^3] As a special case, if we choose a contour $\mathcal{C}$ along the real axis, (5) reduces to the Zel’dovich regularization Zel’dovich (1961) (see also earlier work Kapur et al. (1938)).. We can analytically show that $\mathcal{A}_{n}^{2}$ satisfies a useful formula Motohashi (2024b)
$B_{n}=\frac{iA_{\rm out}^{2}}{2\omega_{n}\mathcal{A}_{n}^{2}},$ (6)
which allows for an alternative way to calculate the excitation factor $B_{n}$ defined by (2). The equivalence between (6) and (2) is also numerically crosschecked for various specific cases including Schwarzschild and Kerr black holes. The formula (6) tells us that the excitation factor $B_{n}$ becomes large when the norm squared $\mathcal{A}_{n}^{2}$ becomes small, which precisely occurs at the avoided crossing near the EP.
Resonant excitation at avoided crossing. To investigate how QNM frequencies and excitation factors change when two QNM frequencies get close due to the variation of $U$, let us regard the continuous changes as an accumulation of infinitesimal changes $\delta U$. If a mode $\omega_{n}$ is far from degeneracy, it is straightforward to generalize the Rayleigh-Schrödinger-like perturbation theory for resonance state Zel’dovich (1961); Perelomov and Zel’dovich (1998) to QNMs and obtain Motohashi (2024b)
$\delta\omega_{n}=-\frac{(\tilde{\psi}_{n}|\delta U_{n}|\psi_{n})}{2\omega_{n}(\tilde{\psi}_{n}|\psi_{n})}=iB_{n}\frac{(\tilde{\psi}_{n}|\delta U_{n}|\psi_{n})}{A_{\rm out}^{2}},$ (7)
which reveals another role of the excitation factor $B_{n}$, i.e., defining the sensitivity of $\delta\omega_{n}$ to $\delta U$.
The avoided crossing, where two (or more) eigenvalues are close to degenerate, requires a special care  [^4] Note that the similarity between the avoided crossing and repulsion of QNM eigenvalues was pointed out in Dias et al. (2022a, b); Davey et al. (2024) in the studies of charged black holes. Here we examine QNM frequencies and also excitation factors and derive the resonant excitation.. We employ a simple analysis of bifurcation theory of eigenvalues about EP Heiss (2000); Seyranian and Mailybaev (2003) to capture the essential features of the resonance. We focus on the nonrotating black hole case with $U=\omega^{2}-V(x)$  [^5] The rotating black hole case may be investigated by applying a prescription of auxiliary eigenvalue recently employed in Isobe et al. (2024) in the context of the bulk-edge correspondence. and assume that two QNM overtones are close $\omega_{1}\approx\omega_{2}$ and they are isolated from other QNMs so that we can approximately treat them as a two-level system. Suppose a small change $V\to V+\delta V$ as a part of continuous changes, causing a change of the QNM frequencies $\omega_{i}\to\omega_{\pm}$, which can be obtained as eigenvalues of the Hamiltonian $H_{ij}=(\tilde{\psi}_{i}|H|\psi_{j})/(\mathcal{A}_{i}\mathcal{A}_{j})$:
$\omega^{2}_{\pm}=\mathcal{E}_{c}\pm\sqrt{\mathcal{E}_{d}^{2}+\Delta^{2}},$ (8)
where $\mathcal{E}_{c,d}\coloneqq(\mathcal{E}_{1}\pm\mathcal{E}_{2})/2$, $\mathcal{E}_{i}\coloneqq\omega_{i}^{2}+\delta\omega_{i}^{2}$ with $i=1,2$, and $\Delta\coloneqq(\tilde{\psi}_{1}|\delta V|\psi_{2})/(\mathcal{A}_{1}\mathcal{A}_{2})$. When $\Delta$ is negligible, (8) recovers $\mathcal{E}_{i}$, the result of perturbation theory. On the other hand, $\omega_{\pm}$ are degenerate if $\mathcal{E}_{d}^{2}+\Delta^{2}=0$ is satisfied, which defines EP. Since this condition involves complex variables, it is in general not possible to be satisfied by tuning a single real parameter, resulting the avoided crossing.
Let us consider the case where continuous variation of $V$ is driven by a single parameter. At each point except the avoided crossing the perturbation theory is valid, and hence $\omega_{i}^{2}$ follow $\mathcal{E}_{i}$. Suppose that the perturbation theory predicts the two modes approach on the same line, for which, denoting $\mathcal{E}_{d}=pe^{i\alpha}$ and $\mathcal{E}_{c}=\omega_{\rm EP}^{2}$, $p$ is a real parameter and $\alpha$ and $\omega_{\rm EP}^{2}$ are constant. We also neglect the variation of $\Delta=qe^{i\beta}$. The single-parameter degree of freedom of $V$ is thus represented by $p$. Denoting $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}=x+iy$, it is straightforward to derive $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$, where $X+iY$ is defined by rotation of $x+iy$ by $\pi/4-\alpha$. Unless $\alpha=\beta$, which requires a tuning of an additional parameter, $\omega_{\pm}^{2}-\omega_{\rm EP}^{2}$ follows a hyperbolic avoided crossing, and so does $\omega_{\pm}-\omega_{\rm EP}$ itself, since this occurs at a small region near EP. A simple calculation shows $|\omega_{\pm}^{2}-\omega_{\rm EP}^{2}|\approx({\rm const}+p^{2})^{1/4}$ at the peak when $p\ll 1$, which means $|\omega_{+}-\omega_{-}|^{-1}$ obeys the quarter-power Lorentzian.
On the other hand, QNM eigenstates $\psi_{\pm}$ associated to $\omega_{\pm}$ are given by $\psi_{+}=\psi_{1}+T\psi_{2}$ and $\psi_{-}=-T\psi_{1}+\psi_{2}$, where $T=1/(\delta+\sqrt{\delta^{2}-1})$ is governed by a dimensionless complex parameter $\delta=\sqrt{\mathcal{E}_{d}^{2}/\Delta^{2}+1}$. If the two modes are close and then $\delta\approx 0$ and $\mathcal{A}_{1}^{2}\approx\mathcal{A}_{2}^{2}$ are satisfied, the eigenstates are almost degenerate $\psi_{-}\approx i\psi_{+}$, and the norm squared nearly vanish as $\mathcal{A}^{2}_{\pm}\propto\pm\delta$. These behaviors are nothing but what is expected at EP Özdemir et al. (2019); Wiersig (2020); Parto et al. (2021); Ding et al. (2022). With (6) and (8), we obtain $B_{\pm}\propto\mathcal{A}^{-2}_{\pm}\propto(\omega_{\pm}^{2}-\omega_{\rm EP}^{2})^{-1}$. Hence, the excitation factors are amplified following the inverse of the hyperbola, i.e., the lemniscate of Bernoulli, and the resonance peak obeys the quarter-power Lorentzian.
Conclusions and Discussion. When two QNM frequencies approach on the complex plane, they are anomalously excited. We clarified this resonance phenomenon by developing high-precision numerical analysis and theoretical framework of QNMs in the language of non-Hermitian physics. The resonance of Kerr QNMs not only allows us to decipher a long-standing mystery in black hole physics but also stands as a novel beacon for characterizing black hole spacetime geometry. Deviations from Kerr, whether due to astrophysical effects or correction to general relativity, will manifest in the resonances in QNMs. Our findings thus pave the way for rigorous examinations of black holes and the exploration of new physics in gravity.
Acknowledgments. This work was supported by Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Scientific Research (KAKENHI) Grant No. JP22K03639.
Supplemental material
In the main text we focus on the resonance at avoided crossing in the vicinity of EP of Kerr QNMs. To argue that this phenomenon is a universal feature of QNMs, it is helpful to study a simple toy model away from Kerr QNMs. Indeed, the simplest rectangular barrier potential played an important role in a pioneering work Chandrasekhar and Detweiler (1975) in the early days of QNM research. Along the same line, here we consider a simple toy model
$\left(\frac{d^{2}}{dx^{2}}+\omega^{2}-V(x)\right)\psi=0,$ (9)
with double rectangular barriers:
$V(x)=\begin{cases}V_{0},&0\leq x\leq d,\\ V_{1},&b\leq x\leq b+d,\\ 0,&{\rm otherwise},\end{cases}$ (10)
with $0\leq d\leq b$ and $V_{0}\geq 0$ and $V_{1}\geq 0$. We define the QNM frequencies by requiring the purely right-/left-going boundary condition, i.e., $\psi\to e^{\pm i\omega x}$, at $x>b+d$ and $x<0$, respectively. We can then calculate the excitation factor by (2) or (6). This toy model allows us to examine the intrinsic behavior of QNMs and excitation factors in a simple and transparent manner.
Here, we introduce a hierarchy $V_{1}\ll V_{0}$ and investigate the behavior of QNM frequencies and excitation factors with respect to the variation of $b$. For small values of $b$, the QNM spectrum is almost the same as the unperturbed one with $V_{1}=0$, but as we increase $b$, the QNM frequencies migrate on the complex frequency plane.
The QNMs for a similar setup was investigated in Barausse et al. (2014); Cheung et al. (2022); Kyutoku et al. (2023) as the simplest variation of the “flea on the elephant” effect Jona-Lasinio et al. (1981); Graffi et al. (1984); Simon (1985); Santarsiero and Gori (2019), which has been explored recently in the context of the spectral instability Nollert (1996); Barausse et al. (2014); Daghigh et al. (2020); Jaramillo et al. (2021); Qian et al. (2021); Cheung et al. (2022); Jaramillo et al. (2022); Berti et al. (2022); Kyutoku et al. (2023); Cardoso et al. (2024). It has been known that small deformation of potential at far distance causes a drastic change of QNM frequencies while changes in the time-domain ringdown signal shows up only at late time. However, the behavior of the excitation factors has not been clarified.
Let us classify the QNM frequencies for the double rectangular barriers (10) into two categories: The first sequence $\omega^{(1)}_{n}$ $(n=0,1,2,\cdots)$ denotes modes that are smoothly connected to the unperturbed modes with $V_{1}=0$, whereas the second sequence $\omega^{(2)}_{k}$ $(k=0,1,2,\cdots)$ denotes modes that show up when $V_{1}>0$ and correspond to the resonance frequencies for the effective cavity between two barriers. For small $b$, the first sequence modes follow the logarithmic spiral around the unperturbed modes which is consistent with Leung et al. (1997). By using the perturbation theory formula (7), we obtain
$\omega^{(1)}_{n}\simeq\bar{\omega}_{n}-\frac{i\bar{B}_{n}V_{1}\sin\bar{\omega}_{n}d}{\bar{\omega}_{n}}e^{i\bar{\omega}_{n}(2b+d)},$ (11)
where $\bar{\omega}_{n}$ and $\bar{B}_{n}$ are unperturbed QNM frequencies and excitation factors, respectively. On the other hand, for the second sequence, we can derive a simple analytical expression under $V_{1}\ll V_{0}$ as
$\omega^{(2)}_{k}\simeq\frac{1}{2b}\left[(2k+1)\pi-i\ln\frac{V_{0}}{V_{1}}\right].$ (12)
Thus, as we increase $b$, the second sequence modes accumulate to the origin of the complex frequency plane, during which the second sequence modes (or more generally, migrating modes including modes that were originally belonging to the first sequence) approach to the first sequence modes, which then begin to migrate together with the second sequence modes.
We find that when two QNM frequencies approach each other during the migration, they exhibit the same behavior observed for the Kerr QNMs; The two QNM frequencies repel each other, during which the excitation factors are point-symmetrically amplified. In the left and middle columns of Fig. 4, we depict the trajectories of QNM frequencies and excitation factors with respect to the variation of $b$. In addition to $b$, we control the second parameter $V_{1}$ to see how the avoided crossing and resonance change. The cases with $V_{1}d^{2}=4.30\times 10^{-3}$ and $4.31\times 10^{-3}$ demonstrate sharp resonances, where the trajectories follow the hyperbola and lemniscate. The two cases switch about 90 degrees on the complex plane from each other, which is predicted at the sign flip of the RHS of $X^{2}-Y^{2}=q^{2}\sin 2(\alpha-\beta)$ in the main text. The cases with $V_{1}d^{2}=3.50\times 10^{-3}$ and $5.50\times 10^{-3}$ show mild resonances. The trajectories of the QNM frequencies and excitation factors can be regarded as distorted versions of the hyperbola and lemniscate, respectively. These sharp and mild resonances reproduce what we observed in the Kerr QNM frequencies and excitation factors in the main text.
We also examine the characteristic resonant peak appearing in the difference of the excitation factors and the inverse of the difference of the QNM frequencies. As shown in the right column of Fig. 4, the peaks are well fitted by the quarter-power Lorentzian, same as the Kerr case. For the sharp resonances, the relative difference between any pair of $\Delta B$, $1/\Delta\omega$ (normalized by multiplying a numerical constant), and the quarter-power Lorentzian remains $\mathcal{O}(1)$in the plotted region. The mild resonances are still better fitted by the quarter-power Lorentzian over the Lorentzian itself.
Note also that the resonance phenomenon occurs between the fundamental mode and third overtone of the first sequence. It implies that, while the resonances for Kerr QNMs studied in the main text occur for higher overtones, the resonances involving the fundamental mode and/or lower overtones are in general not prohibited. Indeed, we analyze the double rectangular barriers with various parameter sets and confirm the resonances for various pairs of two modes approaching each other. It suggests that the resonance is a universal feature of QNMs.
Figure 4: Left and middle columns: QNM frequencies (left column) and excitation factors (middle column) for the double rectangular barriers (10). The fundamental mode $\omega^{(1)}_{0}$ (red) and third overtone $\omega^{(1)}_{3}$ (green) of the first sequence are highlighted, while overtones up to the seventh in the first (colored) and second (black) sequences are depicted. We set parameters $V_{0}d^{2}=16$ and $V_{1}d^{2}=(3.50,4.30,4.31,5.50)\times 10^{-3}$ (from first to fourth row) and depict trajectories with respect to the variation of $b/d$ in the range $1\leq b/d\leq 15$ with arrows indicating the direction of increasing $b/d$. The logarithmic spiral (black dotted), hyperbola (black dashed), and lemniscate of Bernoulli (black dashed) approximate the trajectories well. Right column: $|B^{(1)}_{0}-B^{(1)}_{3}|$ (blue solid) and $1/|\omega^{(1)}_{0}-\omega^{(1)}_{3}|$ (orange dashed), the latter of which is multiplied by a numerical constant to match the maxima. The peak can be well fitted by a quarter power of Lorentzian (black dotted) rather than the Lorentzian itself (gray dotted).